Consider the paraboloid z=x2+y2. The plane 2x−2y+z−7=0 cuts the paraboloid, its intersection being a curve. Find "the natural" parametrization of this curve. Hint: The curve which is cut lies above a circle in the xy-plane which you should parametrize as a function of the variable t so that the circle is traversed counterclockwise exactly once as t goes from 0 to 2*pi, and the paramterization starts at the point on the circle with largest x coordinate. Using that as your starting point, give the parametrization of the curve on the surface.

c(t)=(x(t),y(t),z(t)), where

x(t)=

y(t)=

z(t)=
Answer #1

Given plane 24-2y + 3-7=0 =-24+24+7 0 integsects paraboloid = x²+42 so in ny plane above two surfaces integsects when –2x+2ySo parameterization above Girdless lait)= -1+3 cost) yet= 1+3 sinit) | Osts 247 Now put this value in to find a SO 32-26-173